OpenStudy (anonymous):

For the function f(x)=3x^2 at the point (4,48), find the a) slope of the curve, b) an equation of the tangent line, c) an equation of the normal line.

OpenStudy (anonymous):

To find the slope of the curve at a point, we can use calculus to help us.

OpenStudy (anonymous):

The derivative of f(x) is f '(x) = 6x. To find the slope at the point (4, 48), plug the x-coordinate into f '(x).

OpenStudy (anonymous):

f '(4) = 6 * (4) = ...?

OpenStudy (anonymous):

Are you there? @tambam16

OpenStudy (anonymous):

yup listening

OpenStudy (anonymous):

Alright, I would like you to try participating. What is 5 * 4 =...?

OpenStudy (anonymous):

Sorry 6 * 4 = ...?

OpenStudy (anonymous):

so it would be f(x)=6(4) is 24.. and?

OpenStudy (anonymous):

24 is the slope of the of the curve at (4, 48)

OpenStudy (anonymous):

This means the answer to part a is 24.

OpenStudy (anonymous):

oh k

OpenStudy (anonymous):

Alright, are you ready for part b?

OpenStudy (anonymous):

yes yes

OpenStudy (anonymous):

To find the equation of a tangent line we can use the slope we got from part a (24), and the point (4, 48) by plugging them into the "point-slope form" of a line: y - y1 = m(x - x1).

OpenStudy (anonymous):

m = slope, x1 = 4, y1 = 48

OpenStudy (anonymous):

y - 48 = 24(x - 4)

OpenStudy (anonymous):

Can you try simplifying that equation?

OpenStudy (anonymous):

y=24x-48

OpenStudy (anonymous):

That's correct! The answer to part b is y = 24x - 48.

OpenStudy (anonymous):

Ready for part c?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

Okay, before we begin part c, I would like to check your understanding of what "normal" means in math. Are you familiar with the meaning of "normal"?

OpenStudy (anonymous):

umm...not really

OpenStudy (anonymous):

Okay, normal means perpendicular. Are you familiar with that word?

OpenStudy (anonymous):

oh yea. so the slope would be -1/24

OpenStudy (anonymous):

Very good! That is correct. Now use the new slope in the same way we did for part b. Plug the slope and the point into the equation y - y1 = m(x - x1)

OpenStudy (anonymous):

m = -1/24, x1 = 4, y1 = 48

OpenStudy (anonymous):

Can you please try plugging that into the equation?

OpenStudy (anonymous):

ok..this is a weird number but it's y=-1/24x + 289/6

OpenStudy (anonymous):

y - 48 = (-1/24)(x - 4)

OpenStudy (anonymous):

y = (-1/24)x + 1/6 + 288/6

OpenStudy (anonymous):

y = (-1/24)x + 289/6

OpenStudy (anonymous):

You got it correct!

OpenStudy (anonymous):

thanks @cebroski !! you are amazing!

OpenStudy (anonymous):

You are welcome! Have a great rest of your night!